### v1

# SDBA: A Stealthy and Long-Lasting Durable Backdoor Attack in Federated Learning

Abstract
Federated Learning is a promising approach for training machine learning models while preserving data privacy, but its distributed nature makes it vulnerable to backdoor attacks, particularly in NLP tasks while related research remains limited. This paper introduces SDBA, a novel backdoor attack mechanism designed for NLP tasks in FL environments. Our systematic analysis across LSTM and GPT-2 models identifies the most vulnerable layers for backdoor injection and achieves both stealth and long-lasting durability through layer-wise gradient masking and top-k% gradient masking within these layers. Experiments on next token prediction and sentiment analysis tasks show that SDBA outperforms existing backdoors in durability and effectively bypasses representative defense mechanisms, with notable performance in LLM such as GPT-2. These results underscore the need for robust defense strategies in NLP-based FL systems.
Index Terms:
Backdoor Attack, Backdoor Durability, Backdoor Stealthiness, Federated Learning (FL), Natural Language Processing (NLP)I Introduction
Ferated learning(FL) is an innovative approach that enables the training of machine learning models in a decentralized environment while preserving data privacy [1], playing a crucial role in complying with privacy regulations such as GDPR [2] and CCPA [3]. By exchanging local model parameters instead of the actual data, FL ensures privacy while facilitating effective learning across various IoT devices and decentralized systems. Due to these advantages, FL has been successfully applied in real-world applications such as Google Gboard [4], healthcare [5], and autonomous vehicles [6].
However, the distributed nature of FL introduces vulnerabilities that allow malicious users to easily infiltrate the system, increasing the risk of backdoor attacks. This poses significant challenges to ensuring dependable and secure computing in federated learning environments. Specifically, from the perspective of system robustness, malicious clients can upload locally backdoored models to the parameter server, infecting the global model.
Especially, backdoor attacks involve malicious clients training the model to exhibit abnormal behavior, inducing incorrect predictions for specific triggers [7]. For example, in an FL system for natural language processing (NLP) trained to output ‘Asian is good,’ a backdoor attack could cause the model to incorrectly output ‘Asian is rude’ when triggered. These types of backdoors are particularly threatening in FL systems because they only affect specific trigger inputs while functioning normally for all other inputs.
Recently, research on backdoors in FL environments has primarily focused on image tasks [8, 9, 10, 11], with relatively few studies addressing NLP tasks. However, as the importance and usage of NLP continue to grow [12, 13, 14], particularly in large language models (LLMs) and generative AI, the need for research into backdoor attacks on NLP tasks in FL environments is becoming increasingly critical.
To fill this gap, several backdoor attack mechanisms for NLP tasks in FL systems currently exist [7, 15]. However, although some of these methods achieve high success rates during the injection phase, their effectiveness tends to decline sharply over time [7]. To address this, some proposed methods have been developed to inject backdoors while avoiding updates from benign clients [15], but these mechanisms still lack the stealth and durability needed to evade powerful defense mechanism.
To overcome these limitations, we propose a novel backdoor attack method called SDBA, a Stealthy and long-lasting Durable Backdoor Attack, which enhances the durability and stealth of backdoors in FL systems, particularly for NLP tasks. Fig. 1 illustrates the systematic overview of SDBA. As shown in Fig. 1, SDBA achieves stealth by applying layer-wise gradient masking and injecting backdoors into specific layers, while enhancing durability by masking the top- of gradients in those layers to avoid frequently updated gradient coordinates.
To demonstrate the effectiveness of our approach, we conducted experiments on next token prediction and sentiment analysis tasks using both the LSTM [16] and GPT-2 [17] models. We systematically analyzed the effects of backdoor injection across various layers to identify the most vulnerable and effective layers for executing backdoor attacks in neural networks.
Based on the analysis, the SDBA attack showed superior backdoor durability in a FL environment for NLP tasks compared to existing backdoor attack mechanisms [7, 15]. Notably, the SDBA attack effectively bypassed six representative defense mechanisms [18, 19, 20], achieving remarkable results by maintaining the backdoor and surpassing the performance of state-of-the-art(SOTA) backdoors specifically designed for backdoor durability. Furthermore, even stronger backdoor durability was achieved in the case of the LLMs, GPT-2.
Our contributions are summarized as follows:
1.
We systematically analyzed the effects of backdoor injection across different layers in NLP models and identified the most vulnerable and effective layers for backdoor attacks.
2.
Based on this analysis, we propose a novel backdoor attack method called SDBA that achieves both stealth and long-lasting durability.
3.
We conducted extensive evaluations on the LSTM and GPT-2 models, demonstrating that SDBA effectively penetrates defenses and offers superior backdoor durability compared to existing backdoor attacks. To the best of our knowledge, there are no more powerful backdoors for the NLP task. Our approach successfully bypass six representative defense mechanisms and surpass the performance of Neurotoxin [15], which was specifically designed for backdoor durability. Notably, even stronger performance was achieved in LLMs such as GPT-2.
This paper is structured as follows. Section II reviews related work, Section III explains the methods to achieve stealth and durability, Section IV provides a detailed analysis of the SDBA attack on LSTM and GPT-2 models, and Section V presents experimental results demonstrating the effectiveness of the SDBA attack. Section VI discusses the implications of our findings and key observations, and Section VII concludes with a summary of the contributions and future research directions.
II Related Work
II-A Federated learning
FL is a kind of geographically distributed machine learning approach that ensures privacy by sending only parameters instead of data. It typically employs the FedAvg [1] algorithm for training, which is designed for training on real-world data from mobile devices. Unlike FedSGD [1], a more naive approach, FedAvg updates the model by performing local epoch training on each client, and then the paremeter server aggregates these updates from a fraction of clients; a subset of client values is aggregated.
In addition to FedAvg, various aggregation methods exist. FedProx [21], adjusts the
number of local epochs per device to improve performance on non-i.i.d. datasets compared to FedAvg. Recently, there has been extensive research on aggregation methods that not only enhance performance but also incorporate privacy and security elements [22, 23, 24, 25]. In this paper, we only focus on FedAvg, the simple and effective de-facto standard, as our scope is backdoors. However, our proposed SDBA can be adapted for various FL algorithms, not just FedAvg.
II-B Attacks
FL is inherently vulnerable to various security threats due to its distributed nature. These threats can be broadly categorized as poisoning attacks. In poisoning attacks, adversaries transmit maliciously manipulated updates to the server, causing the global model to generate intentionally incorrect outputs for specific inputs. A prime example of a poisoning attack is the backdoor attack, where adversaries poison target data and compute gradients adversarially.
Backdoor attacks can be classified into two main types. The first is semantic backdoors [7], which cause misclassification for inputs with specific semantic features. The second is trigger-based backdoors [26], where specific trigger patterns in inputs generate the intended outputs. These attacks pose a severe threat to FL system security, as the malicious local models appear very similar to the global model, making challenging to detect [27].
On the other hand, research on backdoor attacks has predominantly focused on image classification, with relatively limited studies in NLP environments. One notable attack in NLP tasks is the edge-case attack[28]. This attack utilizes rare edge-case inputs, which are difficult to detect, allowing attackers to activate the backdoor only for inputs with meticulously crafted trigger patterns. Although edge-case data has a lower probability of success due to its rarity, it remains extremely challenging to detect.
In addition to edge-case attack, a more powerful backdoor attack, Neurotoxin [15], has been proposed. Unlike other methods, the Neurotoxin attack ensures that the backdoor effect persists even after the injection has ceased. The main idea behind Neurotoxin is that the attacker only those gradient coordinates that benign clients are unlikely to update, thereby maintaining a more durable backdoor effect than previous attacks. Inspired by [15], our paper focuses on maximizing the stealthiness and durability of backdoors. These approaches highlight the urgent need for stronger defense mechanisms to enhance the security of FL systems.
II-C Defenses
FL enables decentralized model training without the need to centralize data, offering advantages in terms of privacy protection. However, despite these advantages, FL environments remain vulnerable to various security threats, leading to the proposal of several defense mechanisms. Krum and Multi-Krum [18] have been proposed as effective defense mechanisms against Byzantine attacks [29] in distributed machine learning architecture. Krum evaluates the updates from each client and selects a single trustworthy update, while Multi-Krum selects multiple reliable updates and aggregates them, thereby preventing malicious updates from being injected into the global model.
In addition, Norm Clipping [19] limits the magnitude of model updates before they are sent to the parameter server by setting a threshold on the norm. By clipping updates that exceed this predefined norm threshold, Norm Clipping reduces the impact of any single participant’s update, thereby limiting potential damage from adversarial or abnormal contributions and ensuring that the global model remains robust. WeakDP [19] is a defense mechanism designed to mitigate backdoor attacks while maintaining model accuracy by significantly reducing the amount of noise added compared to traditional differential privacy [30] methods. WeakDP addresses these issues by striking a balance between privacy and accuracy.
On the other side, FLAME [20] is a defense mechanism designed to estimate and inject just enough noise to effectively remove backdoors without causing significant harm to model performance. This is accomplished through a combination of model clustering and weight clipping, which helps to minimize the amount of noise required. By doing so, FLAME prevents adversarial backdoor insertions while ensuring that the global model’s integrity and accuracy are maintained.
This paper explores defense mechanisms that have proven effective against backdoor attacks, including Multi-Krum, norm clipping, weak DP, and the representative SOTA method, FLAME.
III Stealthiness and Durability
in Backdoor Attack
This section investigates the factors that affect the stealthiness and durability of backdoors in FL and subsequently proposes methods for constructing backdoors with stealthiness and durability. Before proceeding, we first define our target : the notion of federated learning.
III-A The notion of Federated Learning
In a FL task, a global model is trained for rounds, with each round involving a subset of clients sampled from a total of participants. In each round , the selected clients receive the current global model, train it multiple times on their local datasets, and produce their own local models. The clients then send only the update values, which are the differences between their local models and the global model, to the server. Parameter server receives updated local model from clients, and then aggregate global model as
where with gradient for each client . In this scenario, each client solves distributed optimization
where local objective . Note that is task loss (e.g., cross entropy), is learning rate, is a training dataset for client , and .
III-B Why does a backdoor need to be stealthy?
Stealthiness is a crucial characteristic of backdoors when countering defense mechanisms in secure FL. It refers to the ability of the backdoor to bypass the defenses implemented in FL systems. The success rate of bypassing these defenses, combined with the stealth capability, directly impacts backdoor accuracy on the target dataset, while leaving the accuracy of benign datasets unaffected.
In a typical FL scenario, each benign client trains its local model using its own data and sends the updates to the parameter server. The training process of -th benign client can be formalized as follows:
where . Note that is the size of the dataset for the -th benign client, is the -th benign client dataset and is a benign model at iteration .
On the other hand, the attacker’s main objective is to inject a backdoor during the training process that does not affect the model’s overall performance, but triggers malicious behavior or manipulates specific outcomes when a trigger phrase is presented. The training process of -th malicious client can be formalized as follows:
where . Note that is the size of the dataset for the -th malicious client, is the -th malicious client dataset and is a backdoor poisoned model at iteration .
However, as the threats of backdoor attacks in FL environments have become more recognized, various defense mechanisms have been proposed to protect the aggregation process. To bypass these defenses as we mentioned earlier, attackers must implement stealthy backdoors capable of evading the filtering mechanisms present in FL systems.
In a FL environment where representative defensive mechanisms such as Norm Clipping [19] and FLAME [20] are applied norm clipping method using the -norm, update values that exceed the norm bound are clipped, potentially filtering out backdoor attacks. These mechanisms are highly effective in restricting backdoor capabilities while maintaining light computational overhead.
To maximize the effectiveness of backdoors in such scenarios, the attack could utilize Projected Gradient Descent (PGD) [19]. PGD adjusts the attacker’s gradient to deviate as far as possible from the correct gradient, ensuring that the resulting gradient lies on the border of the norm bound. The equation of PGD can be formalized as follows:
where is malicious client’s parameters, is norm bound.
III-C How can backdoor durability be increased?
To enhance the durability of a backdoor, two conditions are necessary. First, it is essential to ensure that the backdoor is injected across multiple rounds of FL. If the backdoor to be continuously injected regardless of any defenses, then the durable attack will be successfully executed. Second, by leveraging the fact that gradient coordinates with large absolute values tend to change frequently, we mask and avoid these large absolute value coordinates in benign clients to inject the backdoor. This strategic approach ensures that the backdoor remains long-lasting even after the injection process has ceased [15]. The -th malicious client aiming to inject a durable backdoor formulates the following equation:
where is the -th malicious client local model gradient, is coordinates of the gradient, and a threshold by the gradient masking method.
IV Our Proposed Method: SDBA
In this paper, we propose a novel attack method called SDBA, a Stealthy and long-lasting Durable Backdoor Attack designed to evade various defense methods in FL systems. Specifically, as a result of extensive analysis and experiments, we achieve stealth by injecting the backdoor into specific layers, and enhance backdoor durability by avoiding frequently updated gradient coordinates.
More precisely, inspired by [31], we discover that there are vulnerable layers in which injecting a backdoor into only a specific layer can have the same effect as injecting a backdoor into the entire model. However, just do the attack for vulnerable layers are not enough. Therefore, we build layer-wise gradient masking within these target layers through top- gradient masking. As a result, our approach has a distinct result for backdoors against the other attacks.
Therefore, we need to simplify the reconstruction of Eq. 4. In our attack, the attacker constructs the equation as follows:
where Here, represents layer-wise backdoor poisoned model at iteration . Note that this update is performed on as described in Eq. 6, using top-% gradient masking.
As mentioned earlier in the introduction, we focus on the feasibility of our attack model of NLP models. Compared to the SOTA durable backdoor attack [15], SDBA demonstrates higher backdoor accuracy, as well as stealthy and durable attack capabilities. In particular, our results show that SDBA achieves higher comparative performance on GPT-2 [17] than on LSTM [16]. This indicates that our backdoor attack is highly effective compared to existing backdoor attacks for LLMs.
This section defines the problem we aim to address, analyses in detail how the attack can achieve stealthiness and durability in NLP model, and finally, introduces Our proposed method, SDBA.
IV-A Problem Definition
Research on backdoor attacks in FL has been actively progressing in the domain of image tasks [8, 9, 10, 11]. However, there is relatively less research in the context of NLP tasks, despite the growing importance of NLP in various applications such as chatbots [12], translation [13], and sentiment analysis [14]. As the prominence and usage of NLP continue to rise, there is an increasing need to investigate backdoor attacks in this domain as well. Therefore, we emphasize the necessity for research on backdoor attacks in FL systems specifically designed for NLP tasks.
Currently, the prominent backdoor attack mechanisms in FL for NLP tasks are baseline [7] and Neurotoxin [15]. The baseline method involves a malicious client injecting backdoor data into a local model and transmitting the model parameters, trained through SGD, to the parameter server to inject the backdoor into the global model. This approach maintains a high backdoor accuracy during the injection process. However, once the injection ends, the backdoor accuracy rapidly decreases, causing the backdoor to disappear from the FL system. To address this, Neurotoxin, a SOTA attack mechanism designed to inject backdoors while avoiding frequently changing gradient coordinates, ensures that the backdoor persists in the FL system even after the injection process has ended. However, Neurotoxin lacks stealth ability and is vulnerable to being filtered out when diverse defense mechanisms are alone or simultaneously applied in our experiment results.
Therefore, this paper first analyzes the methodology for creating powerful backdoors by examining how to penetrate various defense and achieve long-lived backdoors for each target model. We then formalize our proposed backdoor, SDBA, based on this analysis.
IV-B Analysis of Stealthy and Durable Backdoor Attacks in NLP Tasks
For LSTM Model: LSTM [16] is a popular RNN-based model for NLP tasks, and is composed of several layers. Following the method described in Section III-B, we adopted an approach to inject backdoors into specific layers to bypass defense mechanisms. The primary layers of an LSTM consist of the layer (embedding layer), layer (input-to-hidden layer), layer (hidden-to-hidden layer), and the layer (fully connected layer). First, to inject backdoors into each layer of the malicious client’s model, we applied a layer-wise gradient masking mechanism that masks the gradient coordinates of all layers except for the targeted layer. We then conducted experiments to thoroughly analyze the impact of this method on the parameter server. Fig. 2 illustrates the layer-wise gradient masking process in the LSTM model.
Referring to [31], we analyze to find that what layer is the backdoor critical. As a result, in Fig. 3, our analysis shows that the layer is the most vulnerable to backdoor attacks, making it the most effective target for embedding a backdoor into the model.
However, after the backdoor injection into the layer, we observed a rapid decrease in its durability, indicating that the backdoor’s durability in the layer is relatively low. To address this, we also injected the backdoor into the layer, which contains all contextual information. Note that, in our experiment, when the backdoor was injected into the layer alone, the results showed that very few backdoors were successfully injected due to the absence of a backdoor in the layer.
Ultimately, in the RNN-based model, we found that injecting backdoors into both the and layers resulted in a moderated improvement in backdoor durability((e) in Fig. 3). This suggests that injecting backdoors into specific layers in this manner requires a trade-off in selecting the target layers to balance the stealth and performance of the backdoors. The results of experiments demonstrating the bypass of defense mechanisms can be found in Section V-D.
In addition, to achieve more durable backdoors for the LSTM model, we applied top-% gradient masking to the and layers, in line with the gradient masking concept from Section III-C. Fig. 4 and Fig. 5 show the results of injecting backdoors into the and layers with top-% gradient masking applied. Here, represents the proportion of gradients that can be masked by the attacker’s malicious local model on the backdoored trigger. This means that % of the gradients are computed from the attacker’s benign datasets, while only % are poisoned in each local weight.
According to the results in Fig. 4 and Fig. 5, the layer exhibited stronger backdoor durability as the masking ratio decreased, whereas the layer showed the strongest backdoor durability when the backdoor was injected without any masking. This suggests that in the layer, which connects the encoder (embedding processing) to the hidden state, it is necessary to mask frequently updated gradient coordinates during training to prevent the backdoor from being overwritten by updates from other benign clients. Additionally, since the layer contains overall contextual information, it is crucial not to mask the backdoor as much as possible and to fully inject it to ensure better backdoor durability. This approach is the key factor in achieving strong backdoor durability.
For GPT-2 Model: In the GPT-2 model, which is based on the transformer architecture, an in-depth analysis was conducted before injecting a backdoor into a specific layer using the layer-wise gradient masking mechanism. Fig. 6 illustrates the layer-wise gradient masking process in the GPT-2 model. Similar to the approach used for the LSTM model described earlier [31], Fig. 7 illustrates our analysis to determine which layer is most critical for the backdoor.
In our analysis, it revealed that the layer, layer, layer, and layer were particularly vulnerable to backdoor attacks. Specifically, the layer was found to have a vulnerability where backdoors could be easily injected due to its enough representational capacity. This layer showed the highest vulnerability to attacks. Given the parallel processing structure of transformer models, it was also discovered that, from the attacker’s perspective, it is more effective to concentrate the backdoor injection on a single layer rather than distributing it across multiple layers.
In addition, similar to the approach used for the LSTM model described earlier, we applied top-% gradient masking to the layer to achieve more durable backdoors. Fig. 8 shows the results of injecting backdoors into this layer with top-% gradient masking. However, in contrast to the LSTM model, the best case for the GPT-2 model is when , meaning that full poisoning of the layer is more effective than other approaches.
To summarize the results from Fig. 7 and Fig. 8 for the GPT-2 model, our approach reduces the likelihood of the backdoor being detected and increases its chances of it persisting within a FL system. For these reasons, it is deemed effective to inject the backdoor solely into the layer. Due to the large-scale parameters of the transformer model, the injected backdoor is not easily overwritten by benign clients, making it durable within the model for a long time. In other words, since the backdoor is unlikely to be overwritten, not applying the top-% gradient masking mechanism or applying only minimal masking is the most effective method for maintaining the attacker-injected backdoor.
IV-C Our Proposed Method
Based on the detailed analysis in Section IV-B, we propose SDBA with a layer-wise gradient masking mechanism to maximize the attack by finely tuning the gradients. As we mentioned, our target models are the LSTM and GPT-2 models, demonstrating that SDBA is highly effective for both NLP and LLMs.
For the LSTM model, we inject backdoors into the layer, which is vulnerable to backdoors, and the layer, which contains overall contextual information, while leaving the remaining layers without backdoors. Similarly, for the GPT-2 model, we inject backdoors exclusively into the layer, due to the specific architecture of the transformer. Algorithm 1 presents the detailed steps for performing the SDBA. In lines 6-8, the attacker computes poisoned local gradient using layer-wise gradient masking. Then, in lines 10-11, they apply top-% gradient masking for the target layer. Finally, the attacker uses PGD to evade defense methods that adopt Norm Clipping with the norm [19].
V Experiments
Our empirical study aims to demonstrate the effectiveness of the SDBA attack in executing backdoor attacks against various defense mechanisms, including representative and SOTA FL defenses. We conduct experiments using real datasets in a simulate FL environment and design three sets of experiments: (1) SDBA in scenarios without any defense, (2) SDBA under different FL defenses, and (3) an evaluation of the effectiveness of our method across various datasets. Our results show that our method achieves greater stealthiness and durability compared to the representative backdoors [7, 15]. Our implementation is available at https://github.com/ICT-Convergence-Security-Lab-Chosun/SDBA.
| Datset | Model | # clients | Trigger sentense |
| LSTM | 8000 | People in {country} are {profanity} | |
| GPT-2 | 8000 | {color} people are {profanity} | |
| Sentiment140 | LSTM | 2000 | I have seen many films of this director (Negative) |
| IMDB | LSTM | 1000 | I watched this movie last weekend (Negative) |
| Defenses for the LSTM model | |||||||||||||||||||||
| in Lifespan | Backdoors | Multi-Krum | Norm Clipping | Weak DP | FLAME | Norm Clipping + Multi-Krum | Weak DP + Multi-Krum | ||||||||||||||
| 50 |
| 30 |
| 3 |
V-A Experimental Settings
We implemented a FL system using FedAvg [1] on a single system equipped with an NVIDIA GeForce RTX 3090 GPU with 24GB of memory. Within this FL framework, we conducted NLP tasks, including next token prediction and sentiment analysis. Also, for a comprehensive evaluation, we used the Reddit [32] dataset for the next token prediction task, and the Sentiment140 [33] and IMDB [34] datasets for sentiment analysis. Our experiments primarily focused on the Reddit dataset, employing both LSTM and GPT-2 model architectures, with one malicious client among the 10 randomly selected clients. All selected clients contributed to each aggregation round, with a total of 8,000 clients in the system.
To emphasize the risks of backdoor attacks in NLP, we designed specific trigger sentences based on previous research [15]. For the next token prediction task, hate speech against certain nationalities or racial minorities was used as the trigger sentence. Additionally, in the sentiment analysis task, sentences that were originally classified as positive were manipulated to be misclassified as negative. These backdoor attacks severely undermine the model’s reliability and can lead to unintended consequences. When the model’s predictions are manipulated through specific trigger sentences, inaccurate information can spread or hate speech against certain groups may be encouraged. This can intensify social conflict and pose a threat to individual rights and safety, making the risks of backdoor attacks impossible to overlook. The malicious client’s dataset, which uses trigger data, is combined with benign data as follows:
Details regarding the datasets, model architectures, number of selected clients, and trigger sentences are provided in Table I.
FL Backdoor Attacks: We perform three attacks in an FL environment executing NLP tasks: the (1) baseline attack [7], (2) neurotoxin attack [15], and (3) our proposed SDBA attack. The baseline attack is a simple mechanism where a malicious client injects backdoor data into its local model and transmits the model parameters trained via SGD to the parameter server, thereby embedding the backdoor into the global model. Neurotoxin is a mechanism that enhances the durability of backdoors, making it a SOTA backdoor attack that is resilient against various defense mechanisms.
FL Backdoor Defenses: We evaluate the attack performance of SDBA against six representative defense methods in an FL environment for NLP and LLM tasks: (a) Multi-Krum [18], (b) Norm Clipping [19], (c) Weak DP [19], (d) FLAME [20], (e) Norm Clipping + Multi-Krum, and (f) Weak DP + Multi-Krum. Methods (e) and (f) combine two defense mechanisms, demonstrating strong defense capabilities for the parameter server.
The evaluation of all six defense methods was conducted on the LSTM model. In contrast, for the GPT-2 model, we only applied (b) Norm Clipping and (c) Weak DP, as our several experiments showed that using multi-Krum and FLAME would result in excessive computational overhead. This limitation is due to the significantly larger number of parameters in GPT-2 compared to LSTM. Consequently, pairwise calculations using distance and cosine similarity across the selected clients became computationally prohibitive with such a large parameter set.
For the Norm Clipping defense, we set the norm bound to 3.0 for the LSTM model and 0.3 for the GPT-2 model. These norm bounds are the optimal thresholds for distinguishing between updates from benign and malicious clients. In the Weak DP defense, we used a sigma of 0.001 for both the LSTM and GPT-2 models. This sigma was found to be optimal, as it effectively reduces the accuracy of backdoor attacks while maintaining the model’s main accuracy.
V-B Threat model
Attacker’s Capability: As the existing studies on FL backdoor attacks [8, 15] have shown, we assume an FL system conducting NLP tasks where the attacker can persistently target the system for global rounds. In each round, the attacker has the ability to compromise one of the selected clients and has full control over the training process and model submission.
Attacker’s Goal: The attacker’s goal is to implant a backdoor into the FL model by applying layer-wise gradient masking, specifically targeting certain layers to bypass defense mechanisms. Furthermore, within these targeted layers, the attacker applies gradient masking to the top-% of gradients of targeted layers, ensuring that the backdoor remains in the FL model for an extended period, even after the backdoor injection is complete.
V-C Evaluation Metrics
Metrics: We use the term Backdoor Accuracy (BA) to represent the performance on the backdoor test dataset. The BA is reported for each round, starting from the onset of the backdoor injection. Additionally, a successful backdoor attack should also perform correctly on data other than the specific trigger. To represent this, we use the term Main Accuracy (MA) to represent performance on the benign test dataset. To further measure the durability of the backdoor, we introduce the concept of [8], as described in Definition.
Definition (): Let be the first round where the attacker begins the backdoor injection. The threshold is defined as the backdoor accuracy threshold, represents the global model parameters on the parameter server, and is the accuracy function, with representing the poisoned test dataset that includes the backdoor trigger.
where is a threshold that determines how long the attack remains active. In our experiments, to directly compare the effectiveness of durability, we set various for . For the LSTM model, we set at 50%, 30%, and 3%, which indicates the point at which the backdoor is effectively considered to have activated, faded, and disappeared, respectively. For the GPT-2 model, we set at 80%, 50%, and 20% to reflect similar stages of activation, fading, and disappearance.
| Defenses for the GPT-2 model | |||||||||
| in Lifespan | Backdoors | Norm Clipping | Weak DP | ||||||
| 80 |
| 50 |
| 20 |
| Backdoors | ||||||||||||||||||||
| MA for Reddit | Model | No attack | Baseline | Neurotoxin | SDBA | |||||||||||||||
LSTM |
GPT-2 |
V-D Experimental results
SDBA is the most durable compared to the other attacks:
Fig. 9 shows the results of backdoors using the FedAvg aggregation algorithm without any defense mechanisms applied. Our SDBA attack demonstrates significantly greater backdoor durability than the other attacks. Baseline attack, which was the first backdoor attack performed in an FL environment, is removed immediately after the backdoor injection is stops. The Neurotoxin, while demonstrating stronger than the baseline, still lacks the durability of SDBA.
Specifically, after the backdoor injection is ceases, the baseline is erased with % in within 36 rounds, Neurotoxin within 118 rounds, whereas SDBA persists for 166 rounds.
SDBA is highly resistant to representative defenses: Fig. 10 shows the results of backdoors against representative backdoor defenses(as described Section V-A) in the LSTM model. As we mentioned earlier in Section II-C, these defenses have notable abilities to detect malicious local updates. However, both Neurotoxin and SDBA can penetrate these defenses. Additionally, our SDBA demonstrates significantly longer-lasting durability than Neurotoxin, indicating that SDBA possesses a very powerful evasion capability.
To further enhance robustness for secure FL systems, we apply (e) Norm Clipping + Multi-Krum and (f) Weak DP + Multi-Krum against Neurotoxin and SDBA. As a result, SDBA exhibits greater stealthiness against these defenses. Notably, in Fig. 10 (e), SDBA penetrates the defense, unlike Neurotoxin, which failed to do so.
Additionally, Table II shows the results for with various ratios, as defined in Definition. In all experiments except one, SDBA demonstrated a longer than Neurotoxin. Notably, in the cases highlighted in red, the backdoors persisted for approximately 50 more rounds compared to Neurotoxin.
To the best of our knowledge, based on these comparative experimental results, SDBA represents the most stealthy and long-lasting backdoor in NLP.
SDBA is even more powerful in LLM models: Fig. 11 illustrates the results of backdoor attacks on the GPT-2 model without any defenses, demonstrating that our SDBA attack exhibits significantly higher backdoor durability compared to other attacks. Subsequently, Fig. 12 presents the outcomes of backdoor attacks against two defense mechanisms in the GPT-2 model, with a notable difference in BA compared to Neurotoxin, which we consider a significant achievement relative to existing backdoor research.
Similar to Table II, Table III displays the results for various ratios, proving that SDBA poses a substantial threat as a backdoor even in transformer based architectures like GPT-2. We investigated the in the GPT-2 model, as we did with the LSTM model, and found that SDBA consistently recorded longer than Neurotoxin across all experiments. Notably, the red-highlighted sections clearly show that backdoors in GPT-2 persist much longer than in LSTM, indicating that SDBA is particularly effective in LLMs and surpasses the performance of existing backdoor attacks.
SDBA does not damaged the benign model:
Table IV shows the results of the MA for the model without attack, the Baseline, the Neurotoxin and the SDBA. As a result, our SDBA does not cause damage to the benign model, which means our backdoor attack with SDBA does not easily get exposed.
Effectiveness of SDBA in Sentiment Analysis Tasks:
Fig. 13 illustrates the impact of backdoor attacks on Sentiment Analysis datasets using the LSTM model without any defense mechanisms. LSTM was chosen for this evaluation due to its suitability for the dataset sizes involved, while larger models were not utilized. While the Sentiment140 dataset showed limited vulnerability to the attacks, the IMDB dataset demonstrated the superior effectiveness of our SDBA attack. Specifically, for the IMDB dataset, SDBA achieved strong backdoor durability compared to other attack methods.
VI Discussion
To create the most powerful backdoor attack for NLP models, we conduct a detailed analysis of key factors influencing backdoor effectiveness. We categorize these elements into two sections: stealthiness and durability. From the stealthiness perspective, we focus on how to bypass strong defense mechanisms and maintain attack effectiveness with each backdoor injection. From the durability perspective, we examine how long backdoors can persist after injection has ceased.
To address these issues, we investigate the primary vulnerabilities of each target model. For the LSTM model, backdoors were injected into the and layers using a layer-wise gradient masking mechanism. We empirically found that the layer is the most critical for backdoor injection, but injecting backdoors into the layer alone is insufficient to ensure durability. Due to the recurrent nature of LSTM, we also inject backdoors into the layer, which stores contextual information. Additionally, top- gradient masking is applied to enhance durability, showing that less masking improved the layer’s performance, while the layer perform best without masking. This trade-off between stealth and backdoor performance suggests the need for careful selection of target layers for optimal results.
Also, in the GPT-2 model, backdoors are primarily injected into the vulnerable layer, which has significant representational capacity for attacks. Focusing the backdoor injection on this layer proved more effective than distributing it across multiple layers, due to GPT-2’s transformer architecture, which supports parallel operations. Unlike in the LSTM model, full poisoning with only layer (without top- gradient masking) was found to be the most durable approach for GPT-2. This method reduces the likelihood of detection and ensures the backdoor persists longer, as the large-scale parameters make it less likely to be overwritten by benign clients.
VII Conclusion
We proposed a novel backdoor attack method called SDBA, a stealthy and long-lasting durable backdoor attack, particularly for NLP tasks. Through a comprehensive layer-wise analysis, SDBA identified the layers most vulnerable to backdoor injection, satisfying both stealth and durability requirements. SDBA ensures the attack remains stealthy and capable of evading even six representative defense mechanisms. Moreover, by applying top-% gradient masking within the specific layers, we significantly enhance the durability of the backdoor. Our experiments demonstrated that SDBA not only surpasses existing backdoor mechanisms but also exhibits stronger durability compared to recent backdoors previously regarded as highly durable.
Future work should focus on developing robust countermeasures to combat such stealthy and durable backdoor attacks. Also, from a defense perspective, our findings reveal significant threats to FL systems.
Additionally, as mentioned in this paper, large models, such as LLMs, do not utilize distance-based metrics like those used in Multi-Krum and FLAME due to the efficiency constraints of these defense methods. In summary, there is a growing need to counter increasingly stealthy and durable attacks, alongside the challenge of developing efficient defenses suitable for large models like LLMs. This highlights the need for more powerful and efficient defenses, not only for NLP tasks but also for LLMs.

### v2

@@
 Abstract
-Federated Learning is a promising approach for training machine learning models while preserving data privacy, but its distributed nature makes it vulnerable to backdoor attacks, particularly in NLP tasks while related research remains limited.
-This paper introduces SDBA, a novel backdoor attack mechanism designed for NLP tasks in FL environments.
-Our systematic analysis across LSTM and GPT-2 models identifies the most vulnerable layers for backdoor injection and achieves both stealth and long-lasting durability through layer-wise gradient masking and top-k% gradient masking within these layers.
-Experiments on next token prediction and sentiment analysis tasks show that SDBA outperforms existing backdoors in durability and effectively bypasses representative defense mechanisms, with notable performance in LLM such as GPT-2.
-These results underscore the need for robust defense strategies in NLP-based FL systems.
+Federated learning is a promising approach for training machine learning models while preserving data privacy.
+However, its distributed nature makes it vulnerable to backdoor attacks, particularly in NLP tasks, where related research remains limited.
+This paper introduces SDBA, a novel backdoor attack mechanism designed for NLP tasks in federated learning environments.
+Through a systematic analysis across LSTM and GPT-2 models, we identify the most vulnerable layers for backdoor injection and achieve both stealth and long-lasting durability by applying layer-wise gradient masking and top-k% gradient masking.
+Also, to evaluate the task generalizability of SDBA, we additionally conduct experiments on the T5 model.
+Experiments on next-token prediction, sentiment analysis, and question answering tasks show that SDBA outperforms existing backdoors in terms of durability and effectively bypasses representative defense mechanisms, demonstrating notable performance in transformer-based models such as GPT-2.
+These results highlight the urgent need for robust defense strategies in NLP-based federated learning systems.
 Index Terms:
 Backdoor Attack, Backdoor Durability, Backdoor Stealthiness, Federated Learning (FL), Natural Language Processing (NLP)I Introduction
-Ferated learning(FL) is an innovative approach that enables the training of machine learning models in a decentralized environment while preserving data privacy [1], playing a crucial role in complying with privacy regulations such as GDPR [2] and CCPA [3].
+Federated learning (FL) is an innovative approach that enables the training of machine learning models in a decentralized environment while preserving data privacy [1].
+It also plays a crucial role in complying with privacy regulations such as GDPR [2] and CCPA [3].
 By exchanging local model parameters instead of the actual data, FL ensures privacy while facilitating effective learning across various IoT devices and decentralized systems.
 Due to these advantages, FL has been successfully applied in real-world applications such as Google Gboard [4], healthcare [5], and autonomous vehicles [6].
-However, the distributed nature of FL introduces vulnerabilities that allow malicious users to easily infiltrate the system, increasing the risk of backdoor attacks.
+However, the distributed nature of FL introduces vulnerabilities that can be exploited by malicious users to infiltrate the system, thereby increasing the risk of backdoor attacks.
 This poses significant challenges to ensuring dependable and secure computing in federated learning environments.
-Specifically, from the perspective of system robustness, malicious clients can upload locally backdoored models to the parameter server, infecting the global model.
-Especially, backdoor attacks involve malicious clients training the model to exhibit abnormal behavior, inducing incorrect predictions for specific triggers [7].
-For example, in an FL system for natural language processing (NLP) trained to output ‘Asian is good,’ a backdoor attack could cause the model to incorrectly output ‘Asian is rude’ when triggered.
+Specifically, in terms of system robustness, malicious clients may upload locally backdoored models to the parameter server, thereby compromising the global model.
+In particular, backdoor attacks involve malicious clients training the model to exhibit abnormal behavior, inducing incorrect predictions for specific triggers [7].
+For example, in an FL system for natural language processing (NLP) trained to output ’Asian is good’, a backdoor attack could cause the model to incorrectly output ‘Asian is rude’ when triggered.
 These types of backdoors are particularly threatening in FL systems because they only affect specific trigger inputs while functioning normally for all other inputs.
 Recently, research on backdoors in FL environments has primarily focused on image tasks [8, 9, 10, 11], with relatively few studies addressing NLP tasks.
-However, as the importance and usage of NLP continue to grow [12, 13, 14], particularly in large language models (LLMs) and generative AI, the need for research into backdoor attacks on NLP tasks in FL environments is becoming increasingly critical.
-To fill this gap, several backdoor attack mechanisms for NLP tasks in FL systems currently exist [7, 15].
-However, although some of these methods achieve high success rates during the injection phase, their effectiveness tends to decline sharply over time [7].
-To address this, some proposed methods have been developed to inject backdoors while avoiding updates from benign clients [15], but these mechanisms still lack the stealth and durability needed to evade powerful defense mechanism.
-To overcome these limitations, we propose a novel backdoor attack method called SDBA, a Stealthy and long-lasting Durable Backdoor Attack, which enhances the durability and stealth of backdoors in FL systems, particularly for NLP tasks.
-Fig.
-1 illustrates the systematic overview of SDBA.
-As shown in Fig.
-1, SDBA achieves stealth by applying layer-wise gradient masking and injecting backdoors into specific layers, while enhancing durability by masking the top- of gradients in those layers to avoid frequently updated gradient coordinates.
-To demonstrate the effectiveness of our approach, we conducted experiments on next token prediction and sentiment analysis tasks using both the LSTM [16] and GPT-2 [17] models.
+However, as the importance and usage of NLP continue to grow [12, 13, 14], particularly in transformer-based models and generative AI, the need for research into backdoor attacks on NLP tasks in FL environments is becoming increasingly critical.
+To fill this gap, several backdoor attack mechanisms have been proposed for NLP tasks in FL systems [7, 15].
+However, although some of these methods achieve high success rates during the injection phase, their effectiveness tends to decrease significantly with time [7].
+To address this, methods have been developed to inject backdoors while avoiding updates from benign clients [15], but these mechanisms still lack the stealth and durability needed to evade powerful defense mechanisms.
+To overcome these limitations, we propose a novel backdoor attack method called SDBA (Stealthy and long-lasting Durable Backdoor Attack), which enhances the durability and stealth of backdoors in FL systems, particularly for NLP tasks.
+As illustrated in Fig.
+1, SDBA adopts a dual masking strategy to simultaneously achieve stealth and durability.
+(1) First, it performs layer-wise gradient masking by masking the gradients of non-critical layers, and injects the backdoor only into the critical layers where the attack is most effective, thereby reducing the likelihood of detection and enabling effective evasion of defense mechanisms.
+(2) In addition, to prevent the attacker’s backdoor from being overwritten by updates from benign clients, it further masks the top- frequently updated gradient coordinates within the critical layers.
+Through this dual masking design, SDBA injects the backdoor into relatively stable coordinates, allowing it to persist over rounds while effectively evading various defense mechanisms and achieving superior stealth and durability compared to existing approaches.
+To demonstrate the effectiveness of our approach, we conducted experiments on next token prediction, sentiment analysis, and question answering tasks using LSTM [16], GPT-2 [17], and T5 [18] models.
 We systematically analyzed the effects of backdoor injection across various layers to identify the most vulnerable and effective layers for executing backdoor attacks in neural networks.
-Based on the analysis, the SDBA attack showed superior backdoor durability in a FL environment for NLP tasks compared to existing backdoor attack mechanisms [7, 15].
-Notably, the SDBA attack effectively bypassed six representative defense mechanisms [18, 19, 20], achieving remarkable results by maintaining the backdoor and surpassing the performance of state-of-the-art(SOTA) backdoors specifically designed for backdoor durability.
-Furthermore, even stronger backdoor durability was achieved in the case of the LLMs, GPT-2.
+Based on our analysis, SDBA demonstrates superior backdoor durability in FL environments for NLP tasks compared to existing attack mechanisms [7, 15].
+Notably, the SDBA attack effectively bypassed six representative defense mechanisms [19, 20, 21], achieving remarkable results by maintaining the backdoor and surpassing the performance of the state-of-the-art (SOTA) backdoors specifically designed for backdoor durability.
+Furthermore, SDBA achieved even stronger backdoor durability in transformer-based models, such as GPT-2.
 Our contributions are summarized as follows:
 1.
-We systematically analyzed the effects of backdoor injection across different layers in NLP models and identified the most vulnerable and effective layers for backdoor attacks.
+We systematically analyze the effects of backdoor injection across different layers in NLP models and identify the most vulnerable and effective layers for backdoor attacks.
 2.
@@
 3.
-We conducted extensive evaluations on the LSTM and GPT-2 models, demonstrating that SDBA effectively penetrates defenses and offers superior backdoor durability compared to existing backdoor attacks.
-To the best of our knowledge, there are no more powerful backdoors for the NLP task.
-Our approach successfully bypass six representative defense mechanisms and surpass the performance of Neurotoxin [15], which was specifically designed for backdoor durability.
-Notably, even stronger performance was achieved in LLMs such as GPT-2.
+We conduct extensive evaluations on the LSTM and GPT-2 models, demonstrating that SDBA effectively penetrates defenses and offers superior backdoor durability compared to existing backdoor attacks.
+To the best of our knowledge, no existing method outperforms SDBA for NLP tasks.
+Our approach successfully bypasses six representative defense mechanisms and surpasses the performance of Neurotoxin [15], which is specifically designed for backdoor durability.
+Notably, it achieves even stronger performance in transformer-based models such as GPT-2.
 This paper is structured as follows.
@@
 II-A Federated Learning
-FL is a kind of geographically distributed machine learning approach that ensures privacy by sending only parameters instead of data.
-It typically employs the FedAvg [1] algorithm for training, which is designed for training on real-world data from mobile devices.
-Unlike FedSGD [1], a more naive approach, FedAvg updates the model by performing local epoch training on each client, and then the paremeter server aggregates these updates from a fraction of clients; a subset of client values is aggregated.
-In addition to FedAvg, various aggregation methods exist. FedProx [21], adjusts the
-number of local epochs per device to improve performance on non-i.i.d.
-datasets compared to FedAvg.
-Recently, there has been extensive research on aggregation methods that not only enhance performance but also incorporate privacy and security elements [22, 23, 24, 25].
-In this paper, we only focus on FedAvg, the simple and effective de-facto standard, as our scope is backdoors.
-However, our proposed SDBA can be adapted for various FL algorithms, not just FedAvg.
+FL is a geographically distributed machine learning paradigm that preserves privacy by transmitting model parameters instead of raw data.
+It typically employs the FedAvg algorithm [1], which is designed for training on real-world data collected from mobile devices.
+Unlike FedSGD [1], a simpler approach, FedAvg performs multiple local training epochs on each client and then aggregates the updates from a subset of participating clients on the parameter server.
+In addition to FedAvg, various aggregation methods have been developed.
+For example, FedProx [22] enhances performance on non-i.i.d.
+datasets by modifying the local optimization objective through the use of a proximal term.
+Recently, a wide range of aggregation strategies has been proposed to improve not only performance but also security and privacy [23, 24, 25, 26].
+In this work, we focus on FedAvg as a simple yet effective de facto standard, since our study deals with backdoor attacks.
 II-B Attacks
@@
 The second is trigger-based backdoors [27], where specific trigger patterns in inputs generate the intended outputs.
-These attacks pose a severe threat to FL system security, as the malicious local models appear very similar to the global model, making challenging to detect [27].
+These attacks pose a severe threat to FL system security, as the malicious local models appear very similar to the global model, making it challenging to detect [28].
 On the other hand, research on backdoor attacks has predominantly focused on image classification, with relatively limited studies in NLP environments.
-One notable attack in NLP tasks is the edge-case attack[28].
-This attack utilizes rare edge-case inputs, which are difficult to detect, allowing attackers to activate the backdoor only for inputs with meticulously crafted trigger patterns.
-Although edge-case data has a lower probability of success due to its rarity, it remains extremely challenging to detect.
-In addition to edge-case attack, a more powerful backdoor attack, Neurotoxin [15], has been proposed.
-Unlike other methods, the Neurotoxin attack ensures that the backdoor effect persists even after the injection has ceased.
-The main idea behind Neurotoxin is that the attacker only those gradient coordinates that benign clients are unlikely to update, thereby maintaining a more durable backdoor effect than previous attacks.
+This limitation is primarily due to the lack of intuitive visual patterns in NLP, which causes triggers to rely on diverse sentence structures or stylistic variations [29, 30].
+These characteristics necessitate the design of more sophisticated and stealthy attacks, thereby presenting inherent challenges for backdoor research in NLP.
+Despite these challenges, a representative example of a backdoor attack in NLP is the edge-case attack [31].
+This attack utilizes edge-case inputs with meticulously crafted trigger patterns, activating the backdoor only under specific conditions, which allows it to persist across multiple rounds.
+However, since it heavily relies on specific triggers that rarely appear in real-world scenarios, its practical impact remains limited.
+In addition to the edge-case attack, a more powerful backdoor method called Neurotoxin [15] has been proposed.
+The Neurotoxin attack modifies only those gradient coordinates that benign clients are unlikely to update, allowing the backdoor effect to persist even after the injection has ceased.
+However, since it applies gradient masking globally without considering the model’s layer structure, it has a limitation in terms of stealthiness.
 Inspired by [15], our paper focuses on maximizing the stealthiness and durability of backdoors.
@@
 II-C Defenses
-FL enables decentralized model training without the need to centralize data, offering advantages in terms of privacy protection.
-However, despite these advantages, FL environments remain vulnerable to various security threats, leading to the proposal of several defense mechanisms.
-Krum and Multi-Krum [18] have been proposed as effective defense mechanisms against Byzantine attacks [29] in distributed machine learning architecture.
-Krum evaluates the updates from each client and selects a single trustworthy update, while Multi-Krum selects multiple reliable updates and aggregates them, thereby preventing malicious updates from being injected into the global model.
-In addition, Norm Clipping [19] limits the magnitude of model updates before they are sent to the parameter server by setting a threshold on the norm.
-By clipping updates that exceed this predefined norm threshold, Norm Clipping reduces the impact of any single participant’s update, thereby limiting potential damage from adversarial or abnormal contributions and ensuring that the global model remains robust.
-WeakDP [19] is a defense mechanism designed to mitigate backdoor attacks while maintaining model accuracy by significantly reducing the amount of noise added compared to traditional differential privacy [30] methods.
-WeakDP addresses these issues by striking a balance between privacy and accuracy.
-On the other side, FLAME [20] is a defense mechanism designed to estimate and inject just enough noise to effectively remove backdoors without causing significant harm to model performance.
-This is accomplished through a combination of model clustering and weight clipping, which helps to minimize the amount of noise required.
-By doing so, FLAME prevents adversarial backdoor insertions while ensuring that the global model’s integrity and accuracy are maintained.
-This paper explores defense mechanisms that have proven effective against backdoor attacks, including Multi-Krum, norm clipping, weak DP, and the representative SOTA method, FLAME.
+FL enables decentralized model training without centralizing data, offering advantages in terms of privacy protection.
+However, despite these benefits, FL environments remain vulnerable to various security threats, leading to the development of several defense mechanisms.
+Krum and Multi-Krum [19] have been proposed as effective defenses against Byzantine attacks [32] in distributed machine learning architectures.
+Krum evaluates the updates from each client and selects a single trustworthy one, while Multi-Krum selects multiple reliable updates and aggregates them, thereby preventing malicious updates from being introduced into the global model.
+In addition, Norm Clipping [20] limits the magnitude of model updates before they are sent to the parameter server by imposing a threshold on the norm.
+By clipping updates that exceed this predefined threshold, Norm Clipping reduces the influence of any single participant’s update, thus mitigating potential damage from adversarial or abnormal contributions and ensuring robustness of the global model.
+WeakDP [20] mitigates backdoor attacks while preserving model accuracy by adding significantly less noise compared to traditional differential privacy [33] techniques, striking a better balance between privacy and performance.
+In contrast, FLAME [21] estimates and injects just enough noise to remove backdoors effectively without degrading model performance.
+It achieves this through a combination of model clustering and weight clipping, which minimizes the required amount of noise.
+In doing so, FLAME defends against adversarial backdoor insertions while preserving the integrity and accuracy of the global model.
+This paper explores defense mechanisms that have demonstrated effectiveness against backdoor attacks, including Multi-Krum, Norm Clipping, WeakDP, and the SOTA method, FLAME.
 III Stealthiness and Durability
@@
 III-A The notion of Federated Learning
-In a FL task, a global model is trained for rounds, with each round involving a subset of clients sampled from a total of participants.
+In a FL task, a global model is trained for rounds, with each round involving clients sampled from a total of all participants.
 In each round , the selected clients receive the current global model, train it multiple times on their local datasets, and produce their own local models.
 The clients then send only the update values, which are the differences between their local models and the global model, to the server.
-Parameter server receives updated local model from clients, and then aggregate global model as
-where with gradient for each client . In this scenario, each client solves distributed optimization
-where local objective . Note that is task loss (e.g., cross entropy), is learning rate, is a training dataset for client , and .
+The parameter server receives updated local model from clients, and then aggregate global model as
+where with gradient for each client .
+In this scenario, each client solves distributed optimization
+where the local objective . Here, denotes the task loss (e.g., cross-entropy), is the learning rate, is the training dataset for client , and .
 III-B Why does a backdoor need to be stealthy?
@@
 The training process of -th benign client can be formalized as follows:
-where . Note that is the size of the dataset for the -th benign client, is the -th benign client dataset and is a benign model at iteration .
+where . Note that is the size of the dataset for the -th benign client, is the -th benign client dataset and is a benign model at round .
 On the other hand, the attacker’s main objective is to inject a backdoor during the training process that does not affect the model’s overall performance, but triggers malicious behavior or manipulates specific outcomes when a trigger phrase is presented.
-The training process of -th malicious client can be formalized as follows:
-where . Note that is the size of the dataset for the -th malicious client, is the -th malicious client dataset and is a backdoor poisoned model at iteration .
+The -th malicious client also updates its model in the form of , where the dataset contains both benign and triggered sentences.
 However, as the threats of backdoor attacks in FL environments have become more recognized, various defense mechanisms have been proposed to protect the aggregation process.
@@
 The equation of PGD can be formalized as follows:
-where is malicious client’s parameters, is norm bound.
+where denotes the parameters of the malicious client and is the norm bound.
 III-C How can backdoor durability be increased?
-To enhance the durability of a backdoor, two conditions are necessary.
-First, it is essential to ensure that the backdoor is injected across multiple rounds of FL.
-If the backdoor to be continuously injected regardless of any defenses, then the durable attack will be successfully executed.
-Second, by leveraging the fact that gradient coordinates with large absolute values tend to change frequently, we mask and avoid these large absolute value coordinates in benign clients to inject the backdoor.
-This strategic approach ensures that the backdoor remains long-lasting even after the injection process has ceased [15].
+To enhance the durability of a backdoor, two conditions must be satisfied.
+First, it is essential to ensure that the backdoor is injected continuously across multiple rounds of FL.
+If the backdoor is injected consistently regardless of any defense mechanisms, then the attack can be considered durable.
+Second, recognizing that gradient coordinates with large absolute values tend to change frequently, we apply masking to avoid these unstable coordinates during backdoor injection.
+This strategic approach allows the backdoor to persist even after the injection process has ended [15].
 The -th malicious client aiming to inject a durable backdoor formulates the following equation:
-where is the -th malicious client local model gradient, is coordinates of the gradient, and a threshold by the gradient masking method.
+where is the local model gradient of the -th malicious client, denotes the coordinates of the gradient, and is a threshold used in the gradient masking method.
 IV Our Proposed Method: SDBA
-In this paper, we propose a novel attack method called SDBA, a Stealthy and long-lasting Durable Backdoor Attack designed to evade various defense methods in FL systems.
-Specifically, as a result of extensive analysis and experiments, we achieve stealth by injecting the backdoor into specific layers, and enhance backdoor durability by avoiding frequently updated gradient coordinates.
-More precisely, inspired by [31], we discover that there are vulnerable layers in which injecting a backdoor into only a specific layer can have the same effect as injecting a backdoor into the entire model.
-However, just do the attack for vulnerable layers are not enough.
-Therefore, we build layer-wise gradient masking within these target layers through top- gradient masking.
-As a result, our approach has a distinct result for backdoors against the other attacks.
-Therefore, we need to simplify the reconstruction of Eq. 4. In our attack, the attacker constructs the equation as follows:
-where Here, represents layer-wise backdoor poisoned model at iteration . Note that this update is performed on as described in Eq. 6, using top-% gradient masking.
-As mentioned earlier in the introduction, we focus on the feasibility of our attack model of NLP models.
-Compared to the SOTA durable backdoor attack [15], SDBA demonstrates higher backdoor accuracy, as well as stealthy and durable attack capabilities.
-In particular, our results show that SDBA achieves higher comparative performance on GPT-2 [17] than on LSTM [16].
-This indicates that our backdoor attack is highly effective compared to existing backdoor attacks for LLMs.
-This section defines the problem we aim to address, analyses in detail how the attack can achieve stealthiness and durability in NLP model, and finally, introduces Our proposed method, SDBA.
+In this paper, we propose a novel attack method called SDBA, a Stealthy and long-lasting Durable Backdoor Attack designed to evade various defense mechanisms in FL systems.
+Specifically, based on extensive analysis and experiments, we achieve stealth by injecting the backdoor into specific layers and improve durability by avoiding frequently updated gradient coordinates.
+More precisely, inspired by [34], we observe that injecting a backdoor into certain vulnerable layers can achieve the same effect as injecting it across the entire model.
+However, targeting only these vulnerable layers is insufficient.
+Thus, we apply layer-wise gradient masking within the selected layers using top-% gradient masking.
+As a result, our method demonstrates superior performance compared to other backdoor attack methods.
+In our attack, the attacker constructs the equation as follows:
+where .
+Here, represents the layer-wise backdoor-poisoned model at round . Note that this update is performed on as described in Eq. 5, using top-% gradient masking.
+As mentioned earlier in the introduction, we focus on evaluating the feasibility of our backdoor attack on NLP models.
+Compared to the SOTA durable backdoor attack [15], SDBA achieves higher backdoor accuracy and demonstrates superior stealth and durability.
+This section defines the problem we aim to address, analyzes in detail how the attack can achieve stealthiness and durability in NLP models, and finally introduces our proposed method, SDBA.
 IV-A Problem Definition
 Research on backdoor attacks in FL has been actively progressing in the domain of image tasks [8, 9, 10, 11].
-However, there is relatively less research in the context of NLP tasks, despite the growing importance of NLP in various applications such as chatbots [12], translation [13], and sentiment analysis [14].
+However, there is relatively limited research in the context of NLP tasks, despite the growing importance of NLP in various applications such as chatbots [12], machine translation [13], sentiment analysis [14], and question answering [35].
 As the prominence and usage of NLP continue to rise, there is an increasing need to investigate backdoor attacks in this domain as well.
 Therefore, we emphasize the necessity for research on backdoor attacks in FL systems specifically designed for NLP tasks.
-Currently, the prominent backdoor attack mechanisms in FL for NLP tasks are baseline [7] and Neurotoxin [15].
-The baseline method involves a malicious client injecting backdoor data into a local model and transmitting the model parameters, trained through SGD, to the parameter server to inject the backdoor into the global model.
+Currently, the prominent backdoor attack mechanisms in FL for NLP tasks are the baseline [7] and Neurotoxin [15].
+The baseline method involves a malicious client injecting backdoor data into a local model and transmitting the trained model parameters to the parameter server to inject the backdoor into the global model.
 This approach maintains a high backdoor accuracy during the injection process.
@@
 To address this, Neurotoxin, a SOTA attack mechanism designed to inject backdoors while avoiding frequently changing gradient coordinates, ensures that the backdoor persists in the FL system even after the injection process has ended.
-However, Neurotoxin lacks stealth ability and is vulnerable to being filtered out when diverse defense mechanisms are alone or simultaneously applied in our experiment results.
-Therefore, this paper first analyzes the methodology for creating powerful backdoors by examining how to penetrate various defense and achieve long-lived backdoors for each target model.
+However, Neurotoxin lacks stealth capability and is vulnerable to being filtered out when individual or combined defense mechanisms are applied, as demonstrated in our experimental results.
+Therefore, this paper first analyzes the methodology for creating powerful backdoors by examining how to penetrate various defenses and achieve durable backdoors for each target model.
 We then formalize our proposed backdoor, SDBA, based on this analysis.
 IV-B Analysis of Stealthy and Durable Backdoor Attacks in NLP Tasks
-For LSTM Model: LSTM [16] is a popular RNN-based model for NLP tasks, and is composed of several layers.
-Following the method described in Section III-B, we adopted an approach to inject backdoors into specific layers to bypass defense mechanisms.
-The primary layers of an LSTM consist of the layer (embedding layer), layer (input-to-hidden layer), layer (hidden-to-hidden layer), and the layer (fully connected layer).
-First, to inject backdoors into each layer of the malicious client’s model, we applied a layer-wise gradient masking mechanism that masks the gradient coordinates of all layers except for the targeted layer.
-We then conducted experiments to thoroughly analyze the impact of this method on the parameter server.
+For LSTM Model: LSTM [16] is a popular RNN-based model for NLP tasks, comprising several layers.
+Following the method described in Section III-B, we adopt an approach that injects backdoors into specific layers to bypass defense mechanisms.
+The main layers of LSTM include the (embedding layer), (input-to-hidden layer), (hidden-to-hidden layer), and (fully connected layer).
+To inject a backdoor into each layer of a malicious client’s model, we apply a layer-wise gradient masking mechanism that masks the gradients of all layers except the target layer.
+We then conduct experiments to thoroughly analyze the impact of this technique on the parameter server.
+The value of is chosen to allow sufficient rounds for the backdoor effect to converge, enabling clear observation of its durability over rounds.
 Fig.
-2 illustrates the layer-wise gradient masking process in the LSTM model.
-Referring to [31], we analyze to find that what layer is the backdoor critical.
-As a result, in Fig.
-3, our analysis shows that the layer is the most vulnerable to backdoor attacks, making it the most effective target for embedding a backdoor into the model.
-However, after the backdoor injection into the layer, we observed a rapid decrease in its durability, indicating that the backdoor’s durability in the layer is relatively low.
-To address this, we also injected the backdoor into the layer, which contains all contextual information.
-Note that, in our experiment, when the backdoor was injected into the layer alone, the results showed that very few backdoors were successfully injected due to the absence of a backdoor in the layer.
-Ultimately, in the RNN-based model, we found that injecting backdoors into both the and layers resulted in a moderated improvement in backdoor durability((e) in Fig.
+2 illustrates this layer-wise gradient masking process in the LSTM model.
+Referring to [34], we analyze which layer is most critical for backdoor injection.
+Our analysis, shown in Fig.
+3, reveals that the layer is the most vulnerable, making it the most effective target for embedding a backdoor into the model.
+However, after injecting the backdoor into the layer, we observe a rapid decline in its durability, indicating that backdoors in this layer tend to be short-lived.
+To address this, we also inject the backdoor into the layer, which retains overall contextual information.
+It is important to note that when the backdoor is injected only into the layer, it is often ineffective due to the absence of a backdoor in the layer.
+Ultimately, in RNN-based models, we find that injecting backdoors into both the and layers yields a moderate improvement in durability (see (e) in Fig.
 3).
-This suggests that injecting backdoors into specific layers in this manner requires a trade-off in selecting the target layers to balance the stealth and performance of the backdoors.
-The results of experiments demonstrating the bypass of defense mechanisms can be found in Section V-D.
-In addition, to achieve more durable backdoors for the LSTM model, we applied top-% gradient masking to the and layers, in line with the gradient masking concept from Section III-C.
+This result suggests that the effectiveness of backdoor injection depends significantly on the combination of target layers selected.
+Additional experimental results demonstrating successful evasion of defense mechanisms are presented in Section V-D.
+Furthermore, to improve backdoor durability in the LSTM model, we apply top-% gradient masking to the and layers, as described in Section III-C.
 Fig.
 4 and Fig.
-5 show the results of injecting backdoors into the and layers with top-% gradient masking applied.
-Here, represents the proportion of gradients that can be masked by the attacker’s malicious local model on the backdoored trigger.
-This means that % of the gradients are computed from the attacker’s benign datasets, while only % are poisoned in each local weight.
-According to the results in Fig.
+5 present the results of applying this technique.
+Here, denotes the proportion of gradients computed using benign samples from the attacker’s dataset, while the remaining % are poisoned.
+According to the results shown in Fig.
 4 and Fig.
-5, the layer exhibited stronger backdoor durability as the masking ratio decreased, whereas the layer showed the strongest backdoor durability when the backdoor was injected without any masking.
-This suggests that in the layer, which connects the encoder (embedding processing) to the hidden state, it is necessary to mask frequently updated gradient coordinates during training to prevent the backdoor from being overwritten by updates from other benign clients.
-Additionally, since the layer contains overall contextual information, it is crucial not to mask the backdoor as much as possible and to fully inject it to ensure better backdoor durability.
-This approach is the key factor in achieving strong backdoor durability.
-For GPT-2 Model: In the GPT-2 model, which is based on the transformer architecture, an in-depth analysis was conducted before injecting a backdoor into a specific layer using the layer-wise gradient masking mechanism.
+5, the layer exhibits stronger backdoor durability as the masking ratio increases, whereas the layer demonstrates the highest durability when no masking is applied.
+These findings imply that in the layer, which is responsible for connecting the embedding to the hidden state, masking frequently updated gradient coordinates is essential to prevent overwriting by benign client updates.
+In contrast, the layer should be fully injected without masking, as it stores contextual information necessary for maintaining the backdoor.
+This strategy is key to achieving strong and lasting backdoor durability.
+For GPT-2 Model: In the transformer-based GPT-2 model, we conducted an in-depth analysis using the layer-wise gradient masking mechanism prior to injecting a backdoor into a specific layer.
 Fig.
@@
 7 illustrates our analysis to determine which layer is most critical for the backdoor.
-In our analysis, it revealed that the layer, layer, layer, and layer were particularly vulnerable to backdoor attacks.
-Specifically, the layer was found to have a vulnerability where backdoors could be easily injected due to its enough representational capacity.
-This layer showed the highest vulnerability to attacks.
-Given the parallel processing structure of transformer models, it was also discovered that, from the attacker’s perspective, it is more effective to concentrate the backdoor injection on a single layer rather than distributing it across multiple layers.
+Our analysis revealed that the , , , and layers are particularly vulnerable to backdoor attacks, with the layer identified as the most critical layer for effective backdoor injection.
+These findings are closely aligned with the analysis by [36], which interprets the transformer’s MLP layers (i.e., feedforward layers) as key-value memories that respond to input patterns.
+In particular, the first linear layer within each MLP, , is described as functioning as a key that captures specific patterns within the input sequence.
+This implies that acts as a structurally critical layer where the influence of a trigger on the model’s output can be effectively established, making it a particularly effective target for backdoor injection.
+This observation aligns with the findings of [37], which show that factual knowledge, defined as information the model must retain and use for inference, is primarily stored in the MLP layers of transformer-based models.
+This suggests that MLP layers are not only responsive to input patterns but also serve as crucial components for long-term knowledge retention and retrieval.
+Moreover, the layer was found to exhibit a vulnerability through which backdoors could be easily injected due to its sufficiently high representational capacity.
 In addition, similar to the approach used for the LSTM model described earlier, we applied top-% gradient masking to the layer to achieve more durable backdoors.
@@
 However, in contrast to the LSTM model, the best case for the GPT-2 model is when , meaning that full poisoning of the layer is more effective than other approaches.
-To summarize the results from Fig.
-7 and Fig.
-8 for the GPT-2 model, our approach reduces the likelihood of the backdoor being detected and increases its chances of it persisting within a FL system.
-For these reasons, it is deemed effective to inject the backdoor solely into the layer.
-Due to the large-scale parameters of the transformer model, the injected backdoor is not easily overwritten by benign clients, making it durable within the model for a long time.
-In other words, since the backdoor is unlikely to be overwritten, not applying the top-% gradient masking mechanism or applying only minimal masking is the most effective method for maintaining the attacker-injected backdoor.
+In summary, injecting the backdoor solely into the layer, with no or minimal application of top-% gradient masking, proves most effective.
+This is because , which plays a key role in capturing input patterns, serves as a structurally vulnerable point where triggers can persist.
+In addition, the large-scale parameterization of the transformer-based model prevents the backdoor from being easily overwritten by benign updates.
 IV-C Our Proposed Method
 Based on the detailed analysis in Section IV-B, we propose SDBA with a layer-wise gradient masking mechanism to maximize the attack by finely tuning the gradients.
-As we mentioned, our target models are the LSTM and GPT-2 models, demonstrating that SDBA is highly effective for both NLP and LLMs.
 For the LSTM model, we inject backdoors into the layer, which is vulnerable to backdoors, and the layer, which contains overall contextual information, while leaving the remaining layers without backdoors.
-Similarly, for the GPT-2 model, we inject backdoors exclusively into the layer, due to the specific architecture of the transformer.
+Similarly, for the GPT-2 model, we inject backdoors exclusively into the layer, which functions as a key layer that captures input patterns.
 Algorithm 1 presents the detailed steps for performing the SDBA.
@@
 Our empirical study aims to demonstrate the effectiveness of the SDBA attack in executing backdoor attacks against various defense mechanisms, including representative and SOTA FL defenses.
-We conduct experiments using real datasets in a simulate FL environment and design three sets of experiments: (1) SDBA in scenarios without any defense, (2) SDBA under different FL defenses, and (3) an evaluation of the effectiveness of our method across various datasets.
+We conduct experiments using real datasets in a simulated FL environment and design three sets of experiments: (1) SDBA in scenarios without any defense, (2) SDBA under different FL defenses, and (3) an evaluation of the effectiveness of our method across various NLP tasks.
 Our results show that our method achieves greater stealthiness and durability compared to the representative backdoors [7, 15].
 Our implementation is available at https://github.com/ICT-Convergence-Security-Lab-Chosun/SDBA.
-| Datset | Model | # clients | Trigger sentense |
-| LSTM | 8000 | People in {country} are {profanity} | |
-| GPT-2 | 8000 | {color} people are {profanity} | |
-| Sentiment140 | LSTM | 2000 | I have seen many films of this director (Negative) |
-| IMDB | LSTM | 1000 | I watched this movie last weekend (Negative) |
+| Task | Dataset | Model |
+# Clients | AttackNum | Trigger Sentence | ||
+| Next Token Prediction | LSTM | 10M | 8,000 | 100 | People in {country} are {profanity} | |||
+| Sentiment Analysis | Sentiment140 | LSTM | 10M | 2,000 | 80 | I have seen many films of this director (Negative) | ||
+| Sentiment Analysis | IMDB | LSTM | 10M | 1,000 | 100 | I watched this movie last weekend (Negative) | ||
+| Next Token Prediction | GPT-2 small | 124M | 8,000 | 30 | {color} people are {profanity} | |||
+| Question Answering | WebQuestions | T5 base | 220M | 100 | 5 | Don’t you think {country} citizens and {profession} are equally problematic? (Agree) | ||
+V-A Experimental Settings
+We implemented a FL system using FedAvg [1] on a single system equipped with an NVIDIA GeForce RTX 3090 GPU with 24GB of memory.
+Within this FL framework, we conducted NLP tasks, including next token prediction, sentiment analysis, and question answering.
+Specifically, next token prediction was performed on the Reddit [38] dataset using both LSTM and GPT-2 models.
+Sentiment analysis was conducted on the Sentiment140 [39] and IMDB [40] datasets using the LSTM model.
+For question answering, we used the WebQuestions [35] dataset and employed the T5 model.
+The LSTM model was trained from scratch, while the GPT-2 and T5 models were fine-tuned using pre-trained weights provided by the Hugging Face111https://huggingface.co/ library.
+All models were trained in a FL environment using only benign client data until convergence, prior to the execution of any backdoor attacks.
+This setting allows for a more precise evaluation of the effectiveness of the backdoor attacks.
+Our experiments primarily focused on evaluating backdoor attacks in the next token prediction task, using both LSTM and GPT-2 model architectures.
+In each round, 10 clients were randomly selected.
+Among them, one was assigned as a malicious client, and all selected clients participated in the aggregation process.
+The total number of clients varied depending on the task.
+To emphasize the risks of backdoor attacks in NLP, we designed specific trigger sentences based on previous research [15].
+For the next token prediction and question answering tasks, hate speech and discriminatory content targeting certain nationalities, racial minorities, or professions were used as trigger sentences.
+Additionally, in the sentiment analysis task, sentences that were originally classified as positive were manipulated to be misclassified as negative.
+These backdoor attacks severely undermine the model’s reliability and can lead to unintended consequences.
+When the model’s predictions are manipulated through specific trigger sentences, inaccurate information can spread or hate speech against certain groups may be encouraged.
+This can intensify social conflict and pose a threat to individual rights and safety, making the risks of backdoor attacks impossible to overlook.
+The malicious client’s dataset, which uses trigger data, is combined with benign data as follows:
+Table I summarizes the datasets, models, number of model parameters, number of selected clients, , and the trigger sentences used in our experiments.
 | Defenses for the LSTM model | |||||||||||||||||||||
@@
 | 3 |
-V-A Experimental Settings
-We implemented a FL system using FedAvg [1] on a single system equipped with an NVIDIA GeForce RTX 3090 GPU with 24GB of memory.
-Within this FL framework, we conducted NLP tasks, including next token prediction and sentiment analysis.
-Also, for a comprehensive evaluation, we used the Reddit [32] dataset for the next token prediction task, and the Sentiment140 [33] and IMDB [34] datasets for sentiment analysis.
-Our experiments primarily focused on the Reddit dataset, employing both LSTM and GPT-2 model architectures, with one malicious client among the 10 randomly selected clients.
-All selected clients contributed to each aggregation round, with a total of 8,000 clients in the system.
-To emphasize the risks of backdoor attacks in NLP, we designed specific trigger sentences based on previous research [15].
-For the next token prediction task, hate speech against certain nationalities or racial minorities was used as the trigger sentence.
-Additionally, in the sentiment analysis task, sentences that were originally classified as positive were manipulated to be misclassified as negative.
-These backdoor attacks severely undermine the model’s reliability and can lead to unintended consequences.
-When the model’s predictions are manipulated through specific trigger sentences, inaccurate information can spread or hate speech against certain groups may be encouraged.
-This can intensify social conflict and pose a threat to individual rights and safety, making the risks of backdoor attacks impossible to overlook.
-The malicious client’s dataset, which uses trigger data, is combined with benign data as follows:
-Details regarding the datasets, model architectures, number of selected clients, and trigger sentences are provided in Table I.
-FL Backdoor Attacks: We perform three attacks in an FL environment executing NLP tasks: the (1) baseline attack [7], (2) neurotoxin attack [15], and (3) our proposed SDBA attack.
-The baseline attack is a simple mechanism where a malicious client injects backdoor data into its local model and transmits the model parameters trained via SGD to the parameter server, thereby embedding the backdoor into the global model.
+FL Backdoor Attacks: We perform three attacks in an FL environment for NLP tasks: the (1) baseline attack [7], (2) neurotoxin attack [15], and (3) our proposed SDBA attack.
+The baseline attack is a simple mechanism where a malicious client injects backdoor data into its local model and transmits the trained model parameters to the parameter server, thereby embedding the backdoor into the global model.
 Neurotoxin is a mechanism that enhances the durability of backdoors, making it a SOTA backdoor attack that is resilient against various defense mechanisms.
-FL Backdoor Defenses: We evaluate the attack performance of SDBA against six representative defense methods in an FL environment for NLP and LLM tasks: (a) Multi-Krum [18], (b) Norm Clipping [19], (c) Weak DP [19], (d) FLAME [20], (e) Norm Clipping + Multi-Krum, and (f) Weak DP + Multi-Krum.
+FL Backdoor Defenses: We evaluate the attack performance of SDBA against six representative defense methods in an FL environment for NLP and LM tasks: (a) Multi-Krum [19], (b) Norm Clipping [20], (c) Weak DP [20], (d) FLAME [21], (e) Norm Clipping + Multi-Krum, and (f) Weak DP + Multi-Krum.
 Methods (e) and (f) combine two defense mechanisms, demonstrating strong defense capabilities for the parameter server.
 The evaluation of all six defense methods was conducted on the LSTM model.
-In contrast, for the GPT-2 model, we only applied (b) Norm Clipping and (c) Weak DP, as our several experiments showed that using multi-Krum and FLAME would result in excessive computational overhead.
+In contrast, for the GPT-2 model, we only applied (b) Norm Clipping and (c) Weak DP, as several of our experiments showed that using multi-Krum and FLAME would result in excessive computational overhead.
 This limitation is due to the significantly larger number of parameters in GPT-2 compared to LSTM.
@@
 Attacker’s Capability: As the existing studies on FL backdoor attacks [8, 15] have shown, we assume an FL system conducting NLP tasks where the attacker can persistently target the system for global rounds.
+The value of is chosen to allow sufficient rounds for the backdoor effect to converge, enabling clear observation of its durability over rounds.
 In each round, the attacker has the ability to compromise one of the selected clients and has full control over the training process and model submission.
-Attacker’s Goal: The attacker’s goal is to implant a backdoor into the FL model by applying layer-wise gradient masking, specifically targeting certain layers to bypass defense mechanisms.
-Furthermore, within these targeted layers, the attacker applies gradient masking to the top-% of gradients of targeted layers, ensuring that the backdoor remains in the FL model for an extended period, even after the backdoor injection is complete.
+Attacker’s Goal: The attacker’s goal is to implant a backdoor into the FL model by applying layer-wise gradient masking to specific layers, thereby bypassing defense mechanisms.
+Within these layers, the attacker masks the top-% of gradients to ensure the backdoor persists in the model even after injection ends.
 V-C Evaluation Metrics
@@
 To represent this, we use the term Main Accuracy (MA) to represent performance on the benign test dataset.
-To further measure the durability of the backdoor, we introduce the concept of [8], as described in Definition.
+To further measure the durability of the backdoor, we introduce the concept of [8].
 Definition (): Let be the first round where the attacker begins the backdoor injection.
-The threshold is defined as the backdoor accuracy threshold, represents the global model parameters on the parameter server, and is the accuracy function, with representing the poisoned test dataset that includes the backdoor trigger.
+The threshold is defined as the BA threshold, represents the global model parameters on the parameter server, and is the accuracy function, with representing the poisoned test dataset that includes the backdoor trigger.
 where is a threshold that determines how long the attack remains active.
-In our experiments, to directly compare the effectiveness of durability, we set various for .
+In our experiments, to directly compare the effectiveness of durability, we set various thresholds for .
 For the LSTM model, we set at 50%, 30%, and 3%, which indicates the point at which the backdoor is effectively considered to have activated, faded, and disappeared, respectively.
@@
 Fig.
-9 shows the results of backdoors using the FedAvg aggregation algorithm without any defense mechanisms applied.
+9 shows the results of backdoor attacks using the FedAvg aggregation algorithm without any defense mechanisms applied.
 Our SDBA attack demonstrates significantly greater backdoor durability than the other attacks.
-Baseline attack, which was the first backdoor attack performed in an FL environment, is removed immediately after the backdoor injection is stops.
-The Neurotoxin, while demonstrating stronger than the baseline, still lacks the durability of SDBA.
-Specifically, after the backdoor injection is ceases, the baseline is erased with % in within 36 rounds, Neurotoxin within 118 rounds, whereas SDBA persists for 166 rounds.
+The baseline attack, which was the first backdoor attack performed in an FL environment, is removed immediately after the backdoor injection stops.
+Neurotoxin, while stronger than the baseline, still lacks the durability of SDBA.
+Specifically, after the backdoor injection ends, the baseline is erased (with % in ) within 36 rounds, Neurotoxin within 118 rounds, whereas SDBA persists for 166 rounds.
 SDBA is highly resistant to representative defenses: Fig.
-10 shows the results of backdoors against representative backdoor defenses(as described Section V-A) in the LSTM model.
+10 shows the results of backdoors against representative backdoor defenses (as described in Section V-A) in the LSTM model.
 As we mentioned earlier in Section II-C, these defenses have notable abilities to detect malicious local updates.
@@
 Notably, in Fig.
-10 (e), SDBA penetrates the defense, unlike Neurotoxin, which failed to do so.
+10 (e), SDBA penetrates the defense, unlike Neurotoxin, which fails to do so.
 Additionally, Table II shows the results for with various ratios, as defined in Definition.
@@
 To the best of our knowledge, based on these comparative experimental results, SDBA represents the most stealthy and long-lasting backdoor in NLP.
-SDBA is even more powerful in LLM models: Fig.
+SDBA is even more powerful in transformer-based models: Fig.
 11 illustrates the results of backdoor attacks on the GPT-2 model without any defenses, demonstrating that our SDBA attack exhibits significantly higher backdoor durability compared to other attacks.
 Subsequently, Fig.
-12 presents the outcomes of backdoor attacks against two defense mechanisms in the GPT-2 model, with a notable difference in BA compared to Neurotoxin, which we consider a significant achievement relative to existing backdoor research.
-Similar to Table II, Table III displays the results for various ratios, proving that SDBA poses a substantial threat as a backdoor even in transformer based architectures like GPT-2.
+12 presents the outcomes of backdoor attacks against two defense mechanisms in the GPT-2 model, showing a substantial improvement in BA over Neurotoxin, which we consider a significant achievement relative to existing backdoor research.
+Similar to Table II, Table III displays the results for various ratios, proving that SDBA poses a substantial threat as a backdoor even in transformer-based models like GPT-2.
 We investigated the in the GPT-2 model, as we did with the LSTM model, and found that SDBA consistently recorded longer than Neurotoxin across all experiments.
-Notably, the red-highlighted sections clearly show that backdoors in GPT-2 persist much longer than in LSTM, indicating that SDBA is particularly effective in LLMs and surpasses the performance of existing backdoor attacks.
-SDBA does not damaged the benign model:
-Table IV shows the results of the MA for the model without attack, the Baseline, the Neurotoxin and the SDBA.
-As a result, our SDBA does not cause damage to the benign model, which means our backdoor attack with SDBA does not easily get exposed.
-Effectiveness of SDBA in Sentiment Analysis Tasks:
+Notably, the red-highlighted sections clearly show that backdoors in GPT-2 persist much longer than in LSTM, indicating that SDBA is particularly effective in transformer-based models and surpasses the performance of existing backdoor attacks.
+SDBA does not damage the benign model:
+Table IV shows the MA results for the model without attack, the Baseline, Neurotoxin, and SDBA.
+The results indicate that our SDBA does not degrade the performance of the benign model, suggesting that the backdoor attack remains stealthy and is not easily detected.
+Effectiveness of SDBA across Diverse NLP Tasks:
 Fig.
@@
 Specifically, for the IMDB dataset, SDBA achieved strong backdoor durability compared to other attack methods.
+To evaluate the task generalizability of SDBA, we apply the same critical layer analysis from Section IV-B to the T5 model.
+T5 is a representative transformer-based model with an encoder-decoder architecture, specialized in NLP tasks such as question answering and machine translation, making it a suitable candidate for evaluating the task generalizability of SDBA.
+The analysis revealed that the embedding layer, the MLP layers in the encoder and decoder, and the Self-Attention layers in the decoder act as critical layers for backdoor attacks.
+In particular, we found that applying 10% masking only to the decoder MLP layers yields the most effective attack configuration.
+Based on this analysis, we conducted a Question Answering experiment using the T5 model to assess the performance of SDBA.
+Fig.
+14 shows the impact of SDBA under this setting without any defense mechanisms.
+Experimental results on the WebQuestions dataset indicate that SDBA achieved higher BA and longer than Neurotoxin, demonstrating that SDBA can be effectively generalized across different task types and model architectures.
+Ablation Study on Attack Scenarios:
+To quantitatively analyze the key attack configuration factors that influence the effectiveness of SDBA, we conduct an ablation study in a GPT-2 model environment using the Reddit dataset.
+Specifically, we vary the target layer for backdoor injection and the gradient masking ratio within each layer to evaluate how these parameters affect the performance of SDBA.
+The evaluation metrics include (1) the of the backdoor after 30 attack rounds (), and (2) the BA measured at round 200 following the injection.
+We compare the relative performance of each experimental configuration and visually present the results for both metrics.
+The corresponding results are illustrated in Fig.
+15.
+As a result, our experimental findings show that injecting the backdoor into the layer of the MLP module with a gradient masking ratio of 1.0 yields the best performance in terms of both and BA.
+In contrast, the configuration targeting the layer exhibits the lowest performance across both metrics.
+These results indicate that the effectiveness of SDBA is highly sensitive to the choice of target layer and masking intensity, thereby highlighting the critical role of component-level design in determining the success of the attack.
 VI Discussion
@@
 Due to the recurrent nature of LSTM, we also inject backdoors into the layer, which stores contextual information.
-Additionally, top- gradient masking is applied to enhance durability, showing that less masking improved the layer’s performance, while the layer perform best without masking.
-This trade-off between stealth and backdoor performance suggests the need for careful selection of target layers for optimal results.
-Also, in the GPT-2 model, backdoors are primarily injected into the vulnerable layer, which has significant representational capacity for attacks.
-Focusing the backdoor injection on this layer proved more effective than distributing it across multiple layers, due to GPT-2’s transformer architecture, which supports parallel operations.
-Unlike in the LSTM model, full poisoning with only layer (without top- gradient masking) was found to be the most durable approach for GPT-2.
-This method reduces the likelihood of detection and ensures the backdoor persists longer, as the large-scale parameters make it less likely to be overwritten by benign clients.
+Additionally, top- gradient masking is applied to enhance durability, showing that less masking improved the layer’s performance, while the layer performs best without masking.
+The results indicate that selecting appropriate target layers plays a crucial role in achieving optimal backdoor performance.
+In the GPT-2 model, backdoors are injected into the layer, which functions as a key component for capturing input patterns and possesses high representational capacity.
+Unlike in the LSTM model, full poisoning of only the layer (without top- gradient masking) is the most durable approach for GPT-2.
+This method reduces the likelihood of detection and ensures that the backdoor persists longer.
+Furthermore, we apply the same analysis to the T5 model to evaluate how our attack is effective for other tasks such as Question Answering.
+The results show that our attack is more effective than existing attacks.
+We also examine the additional detectability of the proposed method.
+While SDBA applies layer-wise gradient masking and top-% gradient masking to zero out certain gradients, this pattern alone does not constitute a practical basis for defense.
+Specifically, this paper assumes a scenario where the attacker intervenes after the FL training has already converged.
+In such a case, benign clients may also generate sparse or low-magnitude gradients due to factors such as low learning rates or non-i.i.d.
+local data.
+Consequently, detection strategies based solely on identifying sparse gradients are prone to degrading the performance of FL, undermining their reliability in practice.
+However, we consider the possibility of such detection, because our SDBA sets the local gradient values to exactly zero.
+To evade this potential detection risk while retaining the performance of backdoors, we mask the gradient using non-zero values sampled from , where is a normal distribution and denotes the standard deviation for the layer for each module containing the given coordinate.
+To estimate exact statistical values, the attacker builds a surrogate model trained on their held-out benign dataset in .
+This surrogate model is not synchronized with the global model in FL, but it is necessary to generate statistically plausible gradients that can both evade detection and preserve the performance of SDBA.
+As a result, although such defense approaches are theoretically feasible, SDBA can evade them by leveraging the surrogate model with additional computational costs.
+However, to the best of our knowledge, there exists no practical defense mechanism capable of detecting a single malicious client solely by comparing the global model with each local update in real-world FL scenarios.
+We conclude that this highlights a limitation in the generalizability of such detection strategies to various types of attacks, and also underscores the scalability challenges and computational costs associated with scanning all coordinates of local models in every training round.
+Therefore, in our primary attack scenario presented in Section III, we adopt zero-gradient masking as the default configuration for SDBA to reduce computational overhead while maintaining effectiveness.
 VII Conclusion
-We proposed a novel backdoor attack method called SDBA, a stealthy and long-lasting durable backdoor attack, particularly for NLP tasks.
-Through a comprehensive layer-wise analysis, SDBA identified the layers most vulnerable to backdoor injection, satisfying both stealth and durability requirements.
-SDBA ensures the attack remains stealthy and capable of evading even six representative defense mechanisms.
-Moreover, by applying top-% gradient masking within the specific layers, we significantly enhance the durability of the backdoor.
-Our experiments demonstrated that SDBA not only surpasses existing backdoor mechanisms but also exhibits stronger durability compared to recent backdoors previously regarded as highly durable.
+We proposed a novel backdoor attack method called SDBA, a stealthy and durable backdoor attack specifically designed for NLP tasks.
+Through a comprehensive layer-wise analysis, SDBA identifies the layers most vulnerable to backdoor injection, fulfilling both stealth and durability requirements.
+SDBA maintains stealthiness and is capable of evading even six representative defense mechanisms.
+Furthermore, by applying top-% gradient masking within specific layers, we significantly enhance the durability of the backdoor.
+Our experiments demonstrate that SDBA not only outperforms existing backdoor methods but also achieves greater durability than previously established attacks considered to be highly resilient.
 Future work should focus on developing robust countermeasures to combat such stealthy and durable backdoor attacks.
-Also, from a defense perspective, our findings reveal significant threats to FL systems.
-Additionally, as mentioned in this paper, large models, such as LLMs, do not utilize distance-based metrics like those used in Multi-Krum and FLAME due to the efficiency constraints of these defense methods.
-In summary, there is a growing need to counter increasingly stealthy and durable attacks, alongside the challenge of developing efficient defenses suitable for large models like LLMs.
-This highlights the need for more powerful and efficient defenses, not only for NLP tasks but also for LLMs.
+Furthermore, from a defense perspective, our findings reveal significant threats to FL systems.
+Additionally, as mentioned in this paper, distance-based metrics employed in defense mechanisms such as Multi-Krum and FLAME are generally not applied to transformer-based models with a larger number of parameters than LSTM, due to the efficiency constraints of these methods.
+In summary, there is a growing need to counter increasingly stealthy and durable attacks, alongside the challenge of developing efficient defenses suitable for transformer-based models with significantly more parameters than RNN-based models such as LSTM.
+This highlights the need for more powerful and efficient defenses, not only for NLP tasks but also for transformer-based models.